### v3

# Vortices in the Many-Body Excited States of Interacting Bosons in Two Dimension

Abstract
Quantum vortices play an important role in the physics of two-dimensional quantum many-body systems, though they usually are understood in the single-particle framework like the mean-field approach. Inspired by the study on the relations between solitons and yrast states, we investigate here the emergence of vortices from the eigenstates of a $N$-body Hamiltonian of interacting particles trapped in a disk. We analyse states that appear by consecutively measuring the positions of particles. These states have densities, phases, and energies closely resembling mean-field vortices. We discuss similarities, but also point out purely many-body features, such as vortex smearing due to the quantum fluctuation of their center. The ab initio analysis of the many-body system, and mean-field approaches are supported by the analysis within the Bogoliubov approach.
preprint: APS/123-QED
Figure 1: Graphical abstract: We study the many-body wave function of interacting bosons on a disk to detect vortices that become apparent after position measurement.
I Introduction
Mean-field (MF) approaches are often the only way to tackle quantum many-body systems and characterise their non-linear effects. At the same time, the interpretation of the MF results in terms of the underlying exact model is sometimes difficult and counter-intuitive. For instance, non-linear effects such as solitons or vortices appear naturally in mean-field frameworks as solutions of dynamical, time-dependent, single-particle, and non-linear mean-field equations. On the other hand, the ab initio model is the linear, many-body Schrödinger equation. Furthermore, it has been shown that the gray solitons, constantly moving non-linear waves known from the MF approach, are related to the non-moving many-body states, which are actually eigenstates of the many-body Hamiltonian [1, 2].
The relation between mean-field solitons and the eigenstates of the many-body Lieb-Liniger model [3, 4], so-called yrast states [5] was a subject of many papers [6, 7, 8, 9, 10, 11, 12, 13, 14]. In between, it has been found that dark soliton, robust dip in a gas density, emerge via breaking symmetry of the system initiated in the yrast state. The symmetry may be broken by simply “measuring” part of the system [7, 8], which in cold gases will happen spontaneously due to particle losses [9]. On the other hand, if the gas is prepared in a product state, with all atoms occupying the mean-field soliton, one finds within Bogoliubov theory quantum correction to this state – the black MF soliton has quantum fluctuations of its position [15] due to anomalous negative mode. The effect of these fluctuations is to gray out the soliton, even though it appears black in the MF description.
The aim of this paper is to put the research further into a study of emergence of 2D vortices out of the many-body yrast states. Quantum vortices, understood via MF approach, play an important role in many contexts, demonstrated in a number of experiments, on rotated superfluid [16], on vortex lattice [17, 18, 19], quantum turbulent gas [20] or the Berezinsky-Kosterlitz transition [21] (see [22] for more information on vortices in ultracold quantum gases). Ongoing research on vortices in 3D turbulence is making use of AI techniques [23, 24, 25] for analyzing numerical [24], and experimental results [25], assuming mean-field description. While mean-field methods are commonly employed in these studies, it is essential to acknowledge their limitations. The observation of unexpected quantum droplets [26], which arise due to the quantum corrections [27], highlights the need to further study model systems at the quantum many-body level and compare them with the MF approach.
In the present work, we focus on a gas of atoms with short-range interactions inside a 2D circular plaquette (see Fig. 1). Unlike in 1D, there are no analytical solutions for vortices – neither in the many-body nor in the MF approach. In fact, even numerically, it is not straightforward to find such dynamical solutions to the MF, that would posses a single vortex with core outside the center of the trap. Our starting points are the many-body yrast states – states minimizing energy at fixed momentum, which we find using the exact diagonalization method with importance truncation [28]. We show that during consecutive “measurements” of positions of particles initiated in an yrast state, one reveals vortices. They may appear off the plaquette center, and at random positions, but are robust dynamically, as confirmed by evolving it with the mean-field equations. In a sense, our complicated procedure serves approximate solution to the MF approach. Special attention is paid to states with angular momentum $j\hbar$, with integer $j$ not larger than the total number of atoms $N$. We show that if $j=N$, the position of the vortex revealed from the quantum states fluctuates. As a result, the average density profile differs between the MF and many-body approach, even in this case. The differences are attributed to the quantum fluctuations, as confirmed by the Bogolubov analysis, similarly to the case of "graying" of the 1D black soliton [15]. Most of the attention is paid to the case $j<N$, in which a vortex emerges out of the center.
In Sec. II we present our model. Sec. III is devoted to the ground state, for which MF gives very good results, for as few as 666 atoms. In Sec. IV we show that vortices are not apparent in the densities, but they emerge in a stochastic process mimicking measurements on many-body excited states. We conclude in Sec V.
II Model
We illustrate our findings for 2D vortices on an example of system of bosons interacting via short range potential and trapped inside a quasi-2 dimensional disk, i.e. assuming infinite potential outside a disk of radius $R$ and neglecting the $Z$ direction. Our Hamiltonian of interest in the second quantized form reads
$\hat{H}$ $=\int\hat{\psi}^{\dagger}({\bm{r}})\left(-\frac{\hbar^{2}}{2M}\nabla^{2}\right)\hat{\psi}({\bm{r}})\mathop{}\!\mathrm{d}{\bm{r}}$ (1)
$+\frac{1}{2}\iint\hat{\psi}^{\dagger}({\bm{r}})\hat{\psi}^{\dagger}({\bm{r}^{\prime}})V({\bm{r}}-{\bm{r}^{\prime}})\hat{\psi}({\bm{r}^{\prime}})\hat{\psi}({\bm{r}})\,\mathop{}\!\mathrm{d}{\bm{r}}\mathop{}\!\mathrm{d}{\bm{r}^{\prime}},$
where $M$ is the atomic mass, $\bm{r}=[x,\,y]$ is a two-dimensional position and $\hat{\psi}(\bm{r})$ is the field operator obeying the bosonic commutation relations $[\hat{\psi}(\bm{r}),\,\hat{\psi}^{\dagger}(\bm{r}^{\prime})]=\delta(\bm{r}-\bm{r}^{\prime})$. Here we model the interaction potential with a repulsive Gaussian $V({\bm{r}})=\frac{g}{\pi\sigma^{2}}\,e^{-\bm{r}^{2}/\sigma^{2}}$, where $g$ is the interaction coupling constant. In what follows we assume that the interaction range $\sigma$ is significantly smaller that the average distance between atoms. The key advantage of using a Gaussian potential, instead of the usual Dirac delta, lies in its convergence – unlike in the Dirac delta potential case, the many-body problem remains well defined without requiring regularization [29] or a suitable adjustment of the interaction strengths [30]. Moreover, for the Gaussian potential, one can evaluate the 2D scattering length analytically to some extent [31].
On the other hand, even for the smooth Gaussian interaction potential, the numerics for $N$ particles are involved (see discussion in [32]). We have achieved satisfactory convergence for $N=6$ atoms using the importance truncation method [28]. Below we discuss without details our numerical approach, whereas the convergence tests are discussed in the B.
Firstly, we decompose the quantum field operator in a single-particle basis as follows:
$\hat{\psi}(r,\phi)=\sum_{j}\psi_{j}(r,\phi)\,\hat{a}_{j},$ (2)
where index $j$ runs over all single-particle states. For our trap geometry those are given as products of radial part and a phase factor
$\psi_{j}(r,\phi)=f_{j}(r)e^{im_{j}\phi},$ (3)
with radial part being proportional to the Bessel function of the first kind
$f_{j}(r)=A_{j}J_{m_{j}}\left(\alpha_{j}\frac{r}{R}\right).$ (4)
For convenience, we label the single-particle eigenstates with a multi-index $j$ that encodes both the magnetic quantum number $m$ and the quantum number labeling the radial eigenstates. The integer $m_{j}$ represents the angular momentum of a given state in units of $\hbar$. For a given $m_{j}$, all radial excitations are described by the first-kind $m_{j}$-th Bessel functions, but rescaled by the factor $\alpha_{j}$. The scaling factors $\alpha_{j}$, which distinguish different radial excitations for a given $m$, are consecutive zeros of the $m_{j}$-th Bessel function. This construction guarantees that the boundary condition $\psi_{j}(R,\phi)=0$ is always satisfied. Finally, $A_{j}$ is the normalization constant. Examples of the radial part of the wavefunctions are given in Fig. 2. Energy of such a single particle state follows the formula
$E_{j}=\frac{\hbar^{2}\alpha_{j}^{2}}{2MR^{2}}.$ (5)
Figure 2: Examples of radial parts of the single particle states used in numerics, Eq. (4) for $m_{j}=0$ (left) and $m_{j}=1$ (right). All radial parts for $m_{j}=0$ are the 0th Bessel functions, but rescaled by $\alpha_{j}$ (with $\alpha_{j}$ being zeros of the 0th bessel function). Similarly, all radial parts for $m_{j}=1$ equal to rescaled 1st Bessel function.
Inserting (2) into (1) we obtain the Hamiltonian in terms of the creation and annihilation operators
$\hat{H}=\sum_{j}E_{j}\hat{a}_{j}^{\dagger}\hat{a}_{j}+\frac{1}{2}\sum_{ijkl}V_{ijkl}\delta^{m_{i}+m_{j}}_{m_{k}+m_{l}}\hat{a}_{i}^{\dagger}\hat{a}_{j}^{\dagger}\hat{a}_{l}\hat{a}_{k},$ (6)
where $V_{ijkl}$ is a double integral
$V_{ijkl}=$ $\int_{0}^{R}\int_{0}^{R}rr^{\prime}f_{i}(r)f_{j}(r^{\prime})f_{k}(r)f_{l}(r^{\prime})$ (7)
×$\times\frac{4\pi g}{\sigma^{2}}e^{-\frac{r^{2}+r^{\prime 2}}{\sigma^{2}}}I_{|m_{i}-m_{k}|}\left(\frac{2rr^{\prime}}{\sigma^{2}}\right)\mathop{}\!\mathrm{d}{r}\mathop{}\!\mathrm{d}{r^{\prime}},$
where $I_{n}(x)$ is the n-th modified Bessel function of the first kind. Note that the Hamiltonian is invariant under rotation, which implies that angular momentum is conserved. As a result, the Kronecker delta appears in Eq. (6). This invariance also implies that all eigenstates of the Hamiltonian can be chosen to be also eigenstates of the angular momentum operator $\hat{L}_{z}$.
Our objects of interest are the excited states of the above Hamiltonian (6), that minimize the energy at a fixed angular momentum, so called yrast states [5]. In our paper we will compare them with the states arising in the simplistic description of the system, the Gross-Pitaevskii equation (GPE), and the Bogoliubov approximations.
Mean-field approximation In the MF approach, it is usually assumed that every particle resides in the same state $\psi_{\mathrm{GP}}(\bm{r},t)$ and its dynamics are governed by the GPE
$i\hbar\frac{\partial}{\partial t}\psi_{\mathrm{GP}}=\left(-\frac{\hbar^{2}}{2M}\nabla^{2}+(N-1)V_{\rm MF}({\bm{r}},t)\right)\psi_{\mathrm{GP}},$ (8)
where the mean-field potential $V_{\rm MF}$ is given by the convolution between the Gaussian interaction potential $V$ and the mean-field density
$V_{\rm MF}({\bm{r}},t)=\int V({\bm{r}}-{\bm{r}}^{\prime})\left|\psi_{\rm GP}({\bm{r}^{\prime}},t)\right|^{2}\mathop{}\!\mathrm{d}{\bm{r}^{\prime}}.$ (9)
We also compare the many-body results with the solutions of the stationary GPE:
$\left(-\frac{\hbar^{2}}{2M}\nabla^{2}+(N-1)V_{\mathrm{MF}}(\bm{r})\right)\psi_{\mathrm{GP}}=\mu\,\psi_{\mathrm{GP}}.$ (10)
Imposing the constrain $\psi_{\mathrm{GP}}(\bm{r})=f(r)e^{i\phi}$ yields a solution with single vortex at the center. In contrast, the ground state can be found assumming $\psi_{\mathrm{GP}}(\bm{r})=f(r)$, which depends only on the distance from the center.
The GPE is the simplest version of the mean-field theory for our system – via non-linear term a single particle “feel” all other particles due to effective potential $V_{\rm MF}$ created by all other particles. We will not invoke in this paper more elaborated mean-field versions, like the ones based on the density-functional theories or generalizations based on the LHY corrections.
Instead we will invoke the Bogoliubov approximation – many-body approach, but based on the assumption that almost all atoms occupy the MF orbital $\psi_{\rm GP}$, with only a small fraction of atoms in other single-particle states [33, 34]. Technically, one has to find the Bogoliubov modes $u(\bm{r})$ and $v(\bm{r})$ that obey the Bogoliubov-de Gennese equations:
$\begin{pmatrix}\hat{H}_{\mathrm{GP}}+\psi_{\mathrm{GP}}\hat{U}^{*}&\psi_{\mathrm{GP}}\hat{U}\\ -\psi_{\mathrm{GP}}^{*}\hat{U}^{*}&-\hat{H}_{\mathrm{GP}}-\psi_{\mathrm{GP}}^{*}\hat{U}\end{pmatrix}\begin{pmatrix}u_{n}\\ v_{n}\end{pmatrix}=\epsilon_{n}\begin{pmatrix}u_{n}\\ v_{n}\end{pmatrix},$ (11)
where
$\hat{H}_{\mathrm{GP}}=-\frac{\hbar^{2}}{2M}\nabla^{2}+(N-1)V_{\mathrm{MF}}(\bm{r})-\mu.$ (12)
We introduce linear operator $\hat{U}$, which depends on $\psi_{\mathrm{GP}}$, and whose action on $u(\bm{r})$ is given by
$\hat{U}u(\bm{r})=(N-1)\int V(\bm{r}-\bm{r}^{\prime})\psi_{\mathrm{GP}}(\bm{r}^{\prime})u(\bm{r}^{\prime})\mathop{}\!\mathrm{d}{\bm{r}^{\prime}}.$ (13)
Solutions can be classified into three families: the “$+$ family”, the “-- family”, and the “0 family”. These names correspond to the normalization of the expression $\langle u_{n}|u_{n}\rangle-\langle v_{n}|v_{n}\rangle$, which equals +1+1+ 1, 1-1- 1 and 0, respectively [33]. Having $u_{n}(\bm{r})$ and $v_{n}(\bm{r})$ one is in position to compute quantities to characterize the system, as the ones listed below.
Quantities In our comparisons between the many-body and GPE approaches, we compute energies and density profiles and then track the differences with the help of the correlation functions.
The energy of the many-body yrast states comes simply from exact diagonalization of the hamiltonian (6). It is compared, wherever possible, with the mean-field energy:
$E_{\rm GP}/N=$ $-\frac{\hbar^{2}}{2M}\int\psi^{*}_{\rm GP}\nabla^{2}\psi_{\rm GP}\mathop{}\!\mathrm{d}{\bm{r}}$ (14)
$+\frac{N-1}{2}$ $\iint\left|\psi_{\rm GP}({\bm{r}})\right|^{2}V({\bm{r}^{\prime}}-{\bm{r}})\left|\psi_{\rm GP}({\bm{r}^{\prime}})\right|^{2}\mathop{}\!\mathrm{d}{\bm{r}^{\prime}}\mathop{}\!\mathrm{d}{\bm{r}},$
and the ground state energy approximated as in the Castin-Dum approach [35] used here to the non-local interaction potential  [^1] The formula for the ground state energy in Bogoliubov (15) was [35] in the context of the short range interaction. We apply it in the same form to our problem.:
$E_{\mathrm{BdG}}=E_{\mathrm{GP}}-\sum_{n\in\mathrm{``+family^{\prime\prime}}}\epsilon_{n}\langle v_{n}|v_{n}\rangle.$ (15)
Another figure of merit is the gas density. In MF the density is simply $N|\psi_{GP}|^{2}$. For many-body system, one has to compute the single particle density matrix (SPDM)
$\rho(\bm{r},\bm{r}^{\prime}):=\langle\Psi|\hat{\psi}^{\dagger}(\bm{r}^{\prime})\hat{\psi}(\bm{r})|\Psi\rangle.$ (16)
Its diagonal, $\rho(\bm{r})=\rho(\bm{r},\bm{r})$, yields information about the average density of particles in the many-body state $|\Psi\rangle$.
The SPDM can be used to test the assumptions underlying MF and BdG approaches, that have the validity range restricted to the weakly interacting limits. Diagonalization of SPDM yields the expansion
$\rho(\bm{r},\bm{r}^{\prime})=N\sum_{i}\lambda_{i}\psi_{i}^{*}(\bm{r}^{\prime})\psi_{i}(\bm{r}),$ (17)
where sum of the coefficients $\lambda_{i}$ is equal to 111. Regime of weak interactions corresponds to situations where majority of the system lies in a single state, say $\psi_{0}$, with its eigenvalue $\lambda_{0}$ being close to 111, that is $\lambda_{0}\gg\lambda_{i}$ for $i\neq 0$.
The value of $\lambda_{0}$ can be compared to the MF fraction $(N-\langle\delta\hat{N}\rangle)/N$ of particles in condensate, where $\langle\delta\hat{N}\rangle$ is the average number of particles out of condensate and can be calculated from BdG solution using formula [33, 34]
$\langle\delta\hat{N}\rangle\simeq\sum_{n\in\mathrm{``+family^{\prime\prime}}}\langle v_{n}|v_{n}\rangle.$ (18)
III Ground state
Figure 3: Energy per atom for states with $N=6$ particles as a function of coupling constant $g$ obtained from the exact diagonalization (solid lines), GPE approach (black dashed lines) and Bogoliubov approach (red dot-dashed lines). The lower curves correspond to the ground state, i.e. 0 total angular momentum, whereas the upper ones to the states minimizing energy with total angular momentum equal to the total number of particles. All results are presented for $\sigma=0.1\ R$.
To ensure that MF and MB (many-body) approaches are mutually consistent, we begin by investigating the ground state. In Fig. 3 we show energy of ground state obtained from exact diagonalization (yellow line) calculated for different coupling constants $g$. In the regime of weak interactions, this result of the exact diagonalization overlaps with results from the mean-field GPE calculations (black line) up to interactions around $gN\approx 5$. The MF energy as proportional to $g$, is just a straight line that gives a correct slope to the exact energy at $g=0$ but it does not capture bending off the energy curve. The latter behaviour is well reproduced by the energy computed within the BdG approach (dot-dashed line) in the whole investigated range of interaction strengths $g$.
To illustrate further the consistency between different approaches we check whether indeed until $gN\approx 5$, the assumption underlying MF and BdG is fulfilled, i.e. if the system is dominated by a single mode. Data for such analysis is illustrated in Fig. 4 showing the relative occupation of the dominant orbital, i.e. $\lambda_{0}$ as defined in Eq. (17). As expected, up to $gN\approx 5$, practically all atoms reside in a single orbital in the ground state.
Figure 4: Condensed fraction of a system described by coefficient $\lambda_{0}$ from equation (17) as a function of the interaction strength. Calculations were performed for $N=6$ particles in the ground state (yellow line) and the yrast state with angular momentum $N\hbar$ (blue line). The red dot-dashed lines illustrate the results obtained from the Bogoliubov approach.
To complete the analysis of the ground state we also compare density profiles, see Fig. 5. It shows the cross-section of MF solution $N|\psi_{\mathrm{GP}}|^{2}$ and the diagonal of SPDM, $\rho(\bm{r})$, for some selected values of the interaction strength $gN$. The density profiles are very similar, even for substantially large interaction strengths.
Figure 5: Ground state density of particles obtained from MF calculations (black dashed) and SPDM from exact diagonalization (yellow solid). Calculations were performed for $N=6$ particles, the interaction strength $gN$ is given in units $\hbar^{2}/M$. Up to the largest considered value $gN=18$, both curves can be seen to overlap. The range of the Gaussian interaction potential equals to $\sigma=0.1\ R$.
To conclude this part – in the weakly interacting regime, i.e. as long as the depletion of the condensate remains small ($\lambda_{0}\approx 1$), the energy and the density profiles are well predicted by the MF approach. In this case, comparing densities is straightforward: the single-particle density obtained from the density matrices of the many-body ground states agrees with the MF densities $N|\psi_{\rm GP}|^{2}$. This indicates a lack of substantial quantum correlations between atoms. In the following, we discuss the yrast states for which the situation is much more subtle.
IV Yrast states
We now turn to the analysis of the eigenstates minimizing energy at a fixed angular momentum, so called yrast states. We will denote the quanta of angular momenta with integer $l$, which will correspond to the value of the angular momentum equal to $l\hbar$. The case of $l=0$ corresponds to the ground state of the system described in the preceding section. We expect that yrast states with $l\neq 0$ contain quantum vortices.
Figure 6: Yrast state density of particles obtained from MF calculations (black dashed) and SPDM from exact diagonalization (blue solid). Calculations were performed for $N=6$ particles and angular momentum $L_{z}=N\hbar$, the interaction strength $gN$ is given in units $\hbar^{2}/M$.
We begin our discussion with the special case $l=N$, where the angular momentum is evenly distributed among all the particles, each carrying $\hbar$. Then, each atom shall rotate with angular momentum $\hbar$ around the disc center, and all atoms altogether will form a single vortex. Similarly to the $l=0$ case, we compare energies (Fig. 3), the quantum depletion (Fig. 4), and the densities (Fig. 6) calculated using the ab initio and MF approaches. Unlike in the ground state case, as the interaction strength increases, differences in results of these two approaches become apparent. The quantum depletion grows much quicker with interaction, as compared with the ground state case. This implies discrepancies in the gas densities. Although the density cross-sections look very similar in both approaches, the MF densities deviates from the SPDM in the center of the system – the density at the center always vanishes in the MF approximation, whereas it is more and more filled with matter in the fully quantum analysis, at least if it comes to the diagonal of a single particle density matrix. We will postpone for a while discussion of our interpretation of the quantum case. The Reader may know similar effect, from the physics of vortices in fermionic superfluid [37] or from the discussion of graying bosonic solitons due to anomalous Bogoliubov mode [15, 38].
More puzzling is the case $0<l<N$. In this case, different scenarios are possible to imagine – off-centered vortex or vortices, with matter rotating around a vortex core, which itself rotates around the center of a trap. It is no straightforward to find numerically such GPE solutions - we haven’t done it. Instead, we study this case starting from the many-body approach – finding yrast state using exact diagonalization method. Our goal here is to show that MF vortices are encoded in the yrast states, and can be revealed via measurements. Firstly, we show the SPDM for $l=N/2$ in Fig. 7. This SPDM does not give a clear view of the system. In particular it shows no vortices. The analysis for the yrast states at a fixed angular momentum $0<l\hbar<N\hbar$ looks as follows: (i) one cannot find easily these states using MF (ii) the density obtained from the single-particle reduced density matrix, Eq. (16), change substantially with interaction and does not show any vortex/vortices. Therefore we use a more subtle approach using high-order correlation function, as described below.
Figure 7: The single-particle density (SPDM) $\rho(\bm{r})$ as a function of a position of a system in an yrast states with angular momentum $L_{z}=N\hbar/2$, obtained via exact diagonalization. There is no clear indication of where the vortex might be located. Parameters: $N=6$ particles, angular momentum $L_{z}=3\hbar$. The interaction strength $gN$ is expressed in units of $\hbar^{2}/M$.
To this end, we use a procedure similar to the one introduced to explain an interference pattern arising in the system in a Fock state [39], also used to find solitons within 1D yrast states [38, 7]. The procedure relies on drawing (“measuring”) positions one-by-one from MB state $|\Psi\rangle$ (see Appendix A for details). One ends up with $N-1$ points $\bar{\bm{r}}_{1},\dots,\bar{\bm{r}}_{N-1}$, which then enter as parameters to the conditional wave function
$\psi_{\textrm{cond}}(\bm{r})\propto\langle\bar{\bm{r}}_{1},\ldots,\bar{\bm{r}}_{N-1},\bm{r}|\Psi\rangle,$ (19)
that is dependent only on one position $\bm{r}$, but it is stochastic. Note that $|\psi_{\textrm{cond}}(\bm{r})|^{2}$ is proportional to the $N$-th order correlation function
$G_{N}=\langle\hat{\psi}^{\dagger}(\bar{\bm{r}}_{1})\ldots\hat{\psi}^{\dagger}(\bar{\bm{r}}_{N-1})\hat{\psi}^{\dagger}(\bm{r})\hat{\psi}(\bm{r})\hat{\psi}(\bar{\bm{r}}_{N-1})\ldots\hat{\psi}(\bar{\bm{r}}_{1})\rangle.$ (20)
We search for vortices within the conditional wavefunctions $\psi_{\mathrm{cond}}$, or in other words – within the high-order correlation function.
Few conditional wave functions for $N=6$ particles, $gN=18\hbar^{2}/M$ and $l=N/2$ are presented in Fig. 8 (a). In most cases one observes a single vortex although its position differs between realisations. To verify whether the vortex-like states emerging in the measuring procedure are related to the MF vortices, we use $\psi_{\rm cond}(\bm{r})$ as initial states for GPE equation and see their dynamics – as in the movies for two interaction strengths shown in the Supplemental Material  [^2] See Supplemental Material.. We found that the vortices are dynamically stable - they rotate around the disc center almost not distorted for weaker interaction $gN=6\hbar^{2}/M$. For stronger interaction $gN=18\hbar^{2}/M$ the system has non-trivial evolution – effect of the beyond mean-field corrections. On other the hand, the density averaged over many measurements, that correspond to the SPDM, will show no vortices – they will be hidden due to the large dispersion in the random position of the vortex core. In the case $l=N$ (Fig. 8 (b)) the dispersion in the vortex position is much smaller, consistent with a clear vortex in the disc center shown in the SPDM in Fig. 6, just slightly greyed-out.
The dispersion of the center of mass around the center of the disk for $l=N$ can also be analyzed using the BdG approach. Numerically, we find the vortex with $l=N$ using the GPE equation and then study its excitations using the BdG equation. Our analysis reveals two modes that contribute most significantly to the quantum depletion, each with roughly equal impact. The first of them corresponds to depletion to $L_{z}=0$ state. Located in the center, it is responsible for filling the vortex core visible in the SPDM in Fig. 6. Its presence is also compatible with the random vortex position in the conditional wave function $\psi_{\mathrm{cond}}$; the distribution of these positions is likely related to the width of this mode. The second mode, with $L_{z}=2\hbar$, ensures the conservation of angular momentum. The densities $|v|^{2}$ of these modes are shown in Fig. 9, along with the vortex profile.
Figure 8: Density and phase of exemplary conditional wave functions (19) for $N=6$ particles, interaction strength $gN=18\hbar^{2}/M$ and the total angular momentum $N\hbar/2=3\hbar$ (a) and $N\hbar=6\hbar$ (b). In both panels there are two columns – the left one for the density and the right one for the phase of $\psi_{\rm cond}$ In both cases, a vortex, if appears, emerges off-centered, while in (a) the variation of its position is much greater.
Figure 9: The cross section of the MF density $N|\psi_{\mathrm{GP}}|^{2}$ (black solid) is compared with the densities $|v_{k}|^{2}$ of two BdG modes that contribute the most to depletion, one with $L_{z}=0$ (green dashed) and one with $L_{z}=2\hbar$ (red dashed). Computations were performed for $N=6$ and $gN=18\hbar^{2}/M$.
V Conclusions
We studied a 2D gas in these excited states that minimize energy at a given angular momentum, known as the yrast states. In a 1D version of this problem, such states are related to mean-field solitons. We have shown that in the 2D version, the states are related to vortices. The vortices emerge spontaneously during measurements, appearing in the high-order correlation functions of the yrast states. When the total angular momentum is less that the number of particles in the system, the system exhibits a single off-center vortex rotating around the center. We demonstrate that the MF-like vortex arising in this manner is always subject to center-of-mass fluctuations, as confirmed by Bogoliubov analysis. The data that support the findings of this article are openly available [41].
VI Acknowledgements
Center for Theoretical Physics of the Polish Academy of Sciences is a member of the National Laboratory of Atomic, Molecular and Optical Physics (KL FAMO).
K.P. acknowledge support from the (Polish) National Science Center Grant No. 2019/34/E/ST2/00289. M.Ś. acknowledge support from the Project no. 2022/45/P/ST3/04237 co-funded by the National Science Centre, Poland and the EU’s H2020 research and innovation programme under the MSCA GA no. 945339.
Appendix A Position Measurement
We describe our procedure of "measuring" $N-1$ positions, one by one, from many-body state $|\Psi\rangle$ to reach the conditional wave-function $\psi_{\rm cond}$ described in Sec. IV. The procedure is recursive starting with $|\Psi_{1}\rangle=|\Psi\rangle$. Measuring the position $\bm{r}_{1}$ of the first particles is done using probability density
$p_{1}(\bm{r})\varpropto\langle\Psi_{1}|\hat{\psi}^{\dagger}(\bm{r})\hat{\psi}(\bm{r})|\Psi_{1}\rangle.$ (21)
Denoting with bar the already drawn position $\bar{\bm{r}}_{1}$, we then calculate (not normalized) state with $N-1$ particles as
$|\Psi_{2}\rangle=\hat{\psi}(\bar{\bm{r}}_{1})|\Psi_{1}\rangle,$ (22)
from which we can draw position of the second particle. Keep in mind that $\bar{\bm{r}}_{1}$ is given from previous, step. In general we draw position of the n-th particle from the distribution
$p_{n}(\bm{r})\varpropto\langle\Psi_{n}|\hat{\psi}^{\dagger}(\bm{r})\hat{\psi}(\bm{r})|\Psi_{n}\rangle,$ (23)
where
$|\Psi_{n}\rangle=\hat{\psi}(\bar{\bm{r}}_{n-1})|\Psi_{n-1}\rangle.$ (24)
Repeating this procedure, we obtain positions of all $N-1$ particles.
Appendix B Numerical methods
In this paper we represent the interatomic potential with a narrow Gaussian of range $\sigma=0.1R$. We study up to 666 atoms – the typical distance between them is ca. $l=\sqrt{\pi R^{2}/6}\approx 0.72R$. As $\sigma$ is substantially smaller than $l$, the interaction potential is practically a short-range one. However, we avoid using the Dirac delta potential directly, as it requires proper regularization in 2D to ensure a finite ground-state energy (see [42] and [43]).
Even with a narrow Gaussian potential, obtaining the ground state necessitates a big numerical basis, making brute-force exact diagonalization impractical. Instead, we employ the importance truncation method (see [28]) to construct an optimal many-body basis for diagonalization. In the importance truncation method one subsequently makes the computational space larger and larger by repeating the so-called pruning and ranking steps. In these steps, one simply evaluates the importance of only those Fock states that appear in the states $\hat{H}|\psi\rangle$, where $|\psi\rangle$ belongs to the computational Hilbert space obtained in the previous steps. This means that for a given cutoff $E_{\rm cut}$, our space consists of only a small fraction of all Fock states with energy smaller than $E_{\rm cut}$. Despite this optimization, our basis still contains up to 300,000 elements for the strongest interaction strengths considered. We then diagonalize the Hamiltonian in this basis using the Lanczos algorithm.
Figure 10 presents a convergence test for the largest interaction strength used, $gN=18\hbar^{2}/M$. The convergence parameter is the energy cutoff $E_{\mathrm{cut}}$, which determines the Fock states included in the importance truncation method. The $E_{\mathrm{cut}}$ is the cutoff for the maximum total kinetic energy of Fock states. Applying a cutoff on the many-body Fock states allows us to access these states more effectively than using a single-particle cutoff.
The left panel shows the ground-state energy as a function of the inverse energy cutoff, while the right panel displays the relative error. For the highest energy cutoff used, the relative error remains below $10^{-5}$, which is sufficient for our qualitative analysis of vortex emergence and quantum effects.
Figure 10: Ground state energy as a function of inverse energy cutoff $E_{\mathrm{cut}}$ (left) and it’s relative error as a function of $E_{\mathrm{cut}}$ (right). The parameter $E_{\mathrm{cut}}$ is the energy cutoff of the whole Fock space. The calculations were performed with importance truncation [28] with parameters $\kappa_{\mathrm{min}}=10^{-5}$ and $C_{\mathrm{cut}}=10^{-4}$.
To conclude, we found that our results are numerically robust.